# Women are less comfortable expressing opinions online than men and report heightened fears for safety: Surveying gender differences in experiences of online harms

### v1

@@ Abstract
Online harms, such as hate speech, trolling and self-harm promotion, continue to be widespread. While some work suggests women are disproportionately affected, other studies find mixed evidence for gender differences in experiences with content of this kind. Using a nationally representative survey of UK adults (N=1992), we examine exposure to a variety of harms, fears surrounding being targeted, the psychological impact of online experiences, the use of safety tools to protect against harm, and comfort with various forms of online participation across men and women. We find that while men and women see harmful content online to a roughly similar extent, women are more at risk than men of being targeted by harms including online misogyny, cyberstalking and cyberflashing. Women are significantly more fearful of being targeted by harms overall, and report greater negative psychological impact as a result of particular experiences. Perhaps in an attempt to mitigate risk, women report higher use of a range of safety tools and less comfort with several forms of online participation, with just 23% of women comfortable expressing political views online compared to 40% of men. We also find direct associations between fears surrounding harms and comfort with online behaviours. For example, fear of being trolled significantly decreases comfort expressing opinions, and fear of being targeted by misogyny significantly decreases comfort sharing photos. Our results are important because with much public discourse happening online, we must ensure all members of society feel safe and able to participate in online spaces.
@@
Online harm , Internet safety , Social media , Online misogyny , Gender-based harm , Safety work , Women’s safety , Public attitudes , Survey research
@@
[label3]Equal contribution as lead authors
@@ 1.1 The problem of online harms
While social media has the potential to bring benefits to its users, such as fostering social connection, helping maintain long distance friendships, facilitating education and empowering people with a platform for their voice, it is also without doubt that social media and the online world more broadly have the potential to expose people to many forms of harm. Harms people may be exposed to online include hate speech, abuse, bullying, misinformation, stalking, image based sexual abuse and self-harm promotion, amongst many others. Common online harms can be categorised as ‘content’ harms (potential online harms that arise from viewing certain posts, pictures, videos, or other content, such as misinformation or eating disorder content), and ‘contact’ harms (potential online harms that originate from direct behaviours towards the target from other users, such as bullying and stalking) (OfCom,, 2021). Recent work has found that 66% of adults in Britain have seen content which they consider to be harmful online, while for younger adults aged 18-34, this is 86% (Enock et al.,, 2023). In this work, we aim to deepen our understanding of people’s key experiences, concerns and behaviours surrounding online harms and online safety with a particular focus on when and how women are disproportionately affected. In particular, we examine the impact of gender, exposure to harm and fears surrounding such exposure on people’s online behaviours in a large, up-to-date nationally representative sample of adults living in the UK.
1.2 Gender differences in experiences of online harms
@@
Some work suggests that women and girls are disproportionately affected by experiences of online harms. For example, online survey research carried out by the UK Victims’ Commissioner office found that for 12 out of 21 categories of online abuse, women were more likely to have been victims than men, with women considerably more likely to experience abuses such as intimate image abuse, cyberstalking and cyberflashing (Storry and Poppleton,, 2022). Additionally, work has found that 90% of those who have been the victim of non-consensual digital distribution of intimate images (also known as revenge porn) are women (United Nations,, 2018). A large-scale review which mapped the state of online violence against women and girls in Europe reported that 9 million girls have experienced some type of online violence by the time they are 15 years old, and across the world, women are 27 times more likely to experience online harassment, leading the authors of the report to describe the internet as a place of gendered violence (European Women’s Lobby,, 2017). Work also suggests that online violence against women often comprises ‘silencing strategies’ such as rape and death threats, attempting to prevent women from partaking in online discussion (Lumsden and Morgan,, 2017). Growing concerns over the problem in recent years has led to the development of the UK Government’s strategy to tackle violence against women and girls in 2021 (Home Office,, 2021), and the recently passed Online Safety Act now requires Ofcom to develop guidance for tech companies to reduce harm to women and girls (End Violence Against Women,, 2023).
@@
However, there is conflicting evidence regarding gender differences in absolute exposure to online harms, with some recent work finding no differences in overall exposure (Enock et al.,, 2023; OfCom,, 2023), and some reports even finding that men experience harms such as ‘sextortion’ (online blackmail where criminals threaten to release sexual/indecent images unless money is paid) and physical threats to a greater extent than women (BBC,, 2023; Pew Research Center,, 2021). One possibility is that much work examining online violence against women in the current literature does not include men in the sample, making it difficult to directly compare gender differences in nuanced experiences of online harms and perhaps emphasising the experiences of women (e.g., Amnesty International,, 2017; European Women’s Lobby,, 2017). Another possibility is that the kinds of harms affecting women online are qualitatively different to those affecting men, with women possibly more at risk of ‘contact’ harms such as sexual harassment and cyberflashing (Levrant Miceli et al.,, 2001; Gillett,, 2023; European Women’s Lobby,, 2017), meaning differences may go undetected when asking about exposure to harms as a whole (e.g., Enock et al.,, 2023). Further, it may be important to consider differences in responses depending on whether people simply say they have seen a harm, compared to whether they have been directly targeted by one. One aim of the present work is to deepen our understanding of the extent to which both men and women are exposed to and targeted by a range of online harms.
@@ 1.3 Psychological impacts of online experiences
As well as considering people’s exposure to online harms and the ways in which individuals may be differentially targeted based on gender, it is also important to consider the psychological impact that exposure to (and even awareness of) various online harms may have. Research has found that exposure to certain harms, such as hate and abuse, can cause severe harm to the psychological wellbeing of targets (Siegel,, 2020). This is also true for ‘harm-advocating content’ such as eating disorder content, self-harm encouragement and sites portraying deaths of others positively, exposure to which was associated with lower subjective wellbeing amongst young Americans and Finns aged 15-30 in a large cross-sectional study of both populations (Keipi et al.,, 2017). Further, a recent systematic review found that victims of cyberstalking and harassment commonly reported experiencing depression, fear, anger, suicidal ideation, shame, paranoia and isolation, amongst other negative emotions (Stevens et al.,, 2021).
If women are targeted by certain online harms to a greater extent than men, it may be that women are disproportionately suffering the negative psychological impact of such exposure. Research by Amnesty found that of the 20% of women who had experienced abuse or harassment through social media, 55% suffered stress, anxiety or panic attacks and 61% had trouble sleeping (Amnesty International,, 2017). Additionally, only 3% of women who had experienced abuse or harassment say they were untroubled by these experiences (compared to 17% of men), while 49% of women (compared to 35% of men) stated that their experiences of online abuse made them feel ashamed (Storry and Poppleton,, 2022). Recent work from OfCom showed that women were less likely than men to say that being online has a positive effect on their mental health (40% of women compared to 45% of men) (OfCom,, 2022). Taken together, there is work to suggest that women may suffer more psychologically as a result of particular online experiences. However, no work to our knowledge directly compares the psychological experiences of men and women in response to online experiences more generally, which is important for knowing where support should most be directed. Another key aim of the present work is to understand whether women are more likely to be negatively affected than men as a direct result of online experiences (for example, through experiencing increased feelings of sadness or anger).
@@ 1.4 Fears surrounding online experiences
Closely tied to negative psychological experience is heightened fear. Recent survey work shows that people’s concerns surrounding the prevalence of online harms is generally high (Bright et al.,, 2024). Based on the fear of crime literature, there is reason to suppose women’s fear of harm overall may be higher than men’s. Offline, women consistently report higher levels of fear of being targeted by crime compared with men, especially when considering crime relating to personal harm (Snedker,, 2012). Stanko, (1995) has argued that women’s anxieties around danger are largely caused by a fear of men, reflecting women’s positions in a gendered world. It is easy to understand why such fears in women may be high. Recent work found that more than one in five young men aged 16 to 29 had a positive view of Andrew Tate, an influencer and self-proclaimed misogynist who has posted content advocating for the hitting and choking of women (Booth,, 2024; Kings College London,, 2024). It is important to understand feelings of fear and safety online as these are likely to impact people’s comfort with participating in the online space and indeed spill over to the offline world too. While some work suggests that women’s fears around personal safety are heightened compared to men’s, no work to date has systematically examined gender differences in fears of specific types of online harm, which is a key aim of the present study.
It is also important to consider where fears online are likely to come from, particularly if there are gender differences in such fears. It is likely, for example, that the media, as a primary source of information for most people, plays an important role in terms of influencing people’s fear of crime (Cashmore,, 2014), particularly owing to the vast coverage of crime-related stories in newspapers and television broadcasts. Little is known about the extent to which perceptions of online harms targeted at friends and public figures, or information from other sources, affects the development of fears surrounding online safety. As a sub-aim, we find out where fears of online harms typically originate from, an important first step in understanding the function of cultural transmission in discussions about online safety.
1.5 Online safety work
1.5.1 Engagement with online safety tools
In considering the impact of exposure to online harms, it is important to understand how, if at all, people attempt to protect themselves from viewing and receiving such content. If women experience greater levels of concern about exposure to online harms, as some prior work suggests, then it is plausible that women engage in more ‘safety work’ in this area to mitigate their risks (Amnesty International,, 2017). Most platforms offer users safety technology such as enabling users to block others, report harmful or inappropriate content, and alter privacy settings on accounts. However, little research has focused on the extent to which people typically engage with such tools, nor the key motivations and demographic drivers underlying their use. Recent survey work from Bright et al., (2024) aimed to address some of these questions, finding that overall awareness and uptake of seven commonly offered safety tools was mixed, with most social media users having heard of tools such as blocking, unfollowing and reporting, and far fewer having heard of altering or deactivating feed algorithms, or hiding engagements with posts. When aware of the tools, women were significantly more likely to use unfollow, block, and report functions than men, and were also more likely to limit who could respond to their content than men, suggesting that women may attempt to do more to protect themselves from harmful or unwanted experiences online. However, the initial work from Bright et al., (2024) asks about use of platform safety tools only amongst users of social media who had heard of the tools, rather than about the use of safety technology amongst the wider population more generally. We aim to examine whether women are more likely than men to engage in safety work online, reflected in increased use of safety tools.
1.5.2 Online participation and safety behaviours
@@
Offline, fear of crime regularly leads to changes in behaviour. As fear of crime is often higher in women, the restrictive behavioural consequences can be greater for women too (Stanko,, 1990; Snedker,, 2012). This can impact women in a myriad of ways, for example in the extra measures they take to ensure they get home safely at night and taking more precautions when travelling alone (Office for National Statistics,, 2022). Just as women have to work harder to protect their safety offline, it looks increasingly likely that this additional burden must be carried into the online world too. Work from Ofcom (2022) found that women are significantly less likely than men to feel that they are able to have a voice online and research exploring academics’ interactions with social media found that female and ethnic minority academics’ fear of harassment online often leads to self-censorship (Olson Carter and LaPoe,, 2018). Taken together, these findings show the silencing effect that online abuse can have, with such harms directly threatening freedom of expression. To deepen our understanding of possible gender differences in comfort with online participation, we examine the extent to which men and women are comfortable engaging in behaviours such as expressing political opinions and sharing content online. Importantly, we test associations between fears about being targeted by online harms and comfort with different forms of online participation.
1.6 Research aims
@@
Despite the existing work discussed so far, we seek to address the gaps in the research landscape that remain surrounding when and how women are disproportionately affected by harmful content online.
@@
Firstly, while much existing work examining online violence against women and girls includes only women and girls in the sample as the key subjects of interest, this means it can be difficult to understand the true extent to which women are affected differently to men. We survey a sample of UK adults nationally representative across age, gender and ethnicity so that we can make meaningful gender comparisons across our variables of interest. Additionally, while much of the existing research either asks about exposure to harmful content in general, or focuses on very specific harms such as abuse and harassment, we ask about exposure and fears relating to a comprehensive list of fifteen potential harms, including content-based ones like misinformation and eating disorder content, along with contact-based ones such as cyberflashing, cyberstalking and image-based sexual abuse. Importantly, we ask not only about general exposure to such harms, but also about whether people have been directly targeted by each one and to what extent. In doing so, we are able to capture a more detailed picture of the different ways in which men and women experience different types of harmful content online.
@@
Our survey design also allows us to deepen our understanding of when and how women are affected differently by online harms to men. Much previous work in the area looks only at how feelings and behaviours change after harm exposure, and only in samples that report having been affected. In our sample, we are able to compare psychological impacts of online experiences between men and women directly, along with fears, the origins of such fears, and subsequent behavioural changes. We are also able to understand whether women engage in more work overall to protect their safety online, including censoring themselves or behaving more cautiously. Importantly, by including a large and heterogeneous sample, we can also examine associations between gender, fears and online participation.
2 Data and Methods
2.1 Data collection, ethics and open science
@@
We conducted a nationally representative survey of 2,000 UK adults. Data collection took place online during June 2023, the survey was created and administered using Qualtrics111www.qualtrics.com and participants were recruited through Prolific [^2] https://www.prolific.com.The survey was approved by the Ethics Committee at The Alan Turing Institute, UK (approval number C2105-074). Informed consent was obtained at the start of the survey according to approved ethical procedures. The materials and data will be available open access on publication.
@@ 2.2 Sample
A total of 1,992 participants who completed the survey passed standard checks for data quality and were included in the final sample. The sample was designed to be nationally representative of the population of the United Kingdom across demographic variables of age, gender and ethnicity using Prolific’s representative sample tool. Respondents were between 18 and 90 years old, with a mean age of 45.8 (SD = 15.5). A total of 1010 participants identified as female (50.7%), 965 as male (48.4%), with 9 as non-binary, 5 selecting ‘prefer not to say’, and 3 self-describing. The majority of respondents identified as White (86.9%), while 7.2% as Asian or Asian British, 3.0% as Black, African, Caribbean, or Black British, and 1.6% as mixed, multiple or other ethnicities (0.7% participants selected ‘any other ethnic group’, while 0.6% chose ‘prefer not to say’) Although participants indicated more specific ethnic identities, we have combined them into broader categories to simplify reporting here. As the main aim of our survey is understanding gender differences, we exclude the 17 participants choosing gender as ’non-binary’, ’prefer not to say’ or ’prefer to self-describe’ from these analyses because the sample size is too small to make meaningful comparisons and inferences for these sub-samples.

### v2

# Gendered Inequalities in Online Harms: Fear, Safety Work, and Online Participation

@@ Abstract
Online harms, such as hate speech, trolling and self-harm promotion, continue to be widespread. There are growing concerns that these harms may disproportionately affect women, reflecting and reproducing existing structural inequalities within digital spaces. Using a nationally representative survey of UK adults (N=1992), we examine how gender shapes exposure to a variety of online harms, fears surrounding being targeted, the psychological impact of online experiences, the use of safety tools, and comfort with various forms of online participation. We find that while men and women report roughly similar levels of absolute exposure to harmful content online, women are more often targeted by contact-based harms including image-based abuse, cyberstalking and cyberflashing. Women report heightened fears about being targeted by online harms, more negative psychological impact in response to online experiences, and increased use of safety tools, reflecting more engagement with personal safety work. Importantly, women also say they are significantly less comfortable with several forms of online participation, for example just 23% of women are comfortable expressing political views online compared to 40% of men. Explanatory models show direct associations between fears surrounding harms and comfort with particular online behaviours. Our findings show how online harms reinforce gender inequality by placing disproportionate psychological burden and participation constraints on women. These results are important because with much public discourse happening online, we must ensure all members of society feel safe and able to participate in online spaces.
@@
Online harms , Internet safety , Social media , Gendered online harms , Safety work , Digital inequality , Public attitudes
note1note1footnotetext: This is a revised version of a manuscript previously posted to arXiv. The prior version remains available for transparency.
@@
[label3]organization=Data Science Institute, The London School of Economics and Political Science, city=London, country=UK
[…]
[label4]organization=Department of Sociology, The London School of Economics and Political Science, city=London, country=UK
@@ 1.1 The problem of online harms
While social media has the potential to bring benefits to its users, such as fostering social connection, facilitating education and empowering people with a platform for their voice, social media and the online world more broadly also have the potential to expose people to many forms of harm. Online harms are typically understood as content or activity that is illegal or that risks causing significant physical or psychological harm to individuals, such as hate speech, abuse, misinformation, image-based sexual abuse and self-harm promotion, among many others. Previous work has found exposure to online harms to be high: 66% of adults in Britain have seen content which they consider to be harmful online, while for younger adults aged 18-34, this is 86% (Enock et al.,, 2023).
1.2 Gendered experiences of online harms
@@
Exposure to online harms can be thought of not only as posing a risk to individuals but also as reflecting and reproducing offline structural inequalities. Building on the concept of the ‘continuum of violence’ posed by Kelly, (1988), feminist scholars have more recently argued that digital abuse and harassment are part of the broader spectrum of gendered violence that extends across offline and online contexts (e.g., Henry and Powell,, 2017; Jane,, 2017). There is some empirical evidence to justify concerns that the impact of online harms reflects this digital extension of existing gendered inequalities, with women and girls often disproportionately affected by these experiences. Online survey research carried out by the UK Victims’ Commissioner office found that for the majority of categories of online abuse, women were more likely to have been victims than men (Storry and Poppleton,, 2022). Work has also found that 90% of those who have been the victim of non-consensual digital distribution of intimate images (sometimes known as ‘revenge porn’) are women (UN Human Rights Council,, 2018). A large-scale review which mapped the state of online violence against women and girls in Europe reported that 9 million girls have experienced some type of online violence by the time they are 15 years old, and across the world, women are 27 times more likely to experience online harassment than men, leading the authors of the report to describe the internet as a place of gendered violence (European Women’s Lobby,, 2017). This concern is corroborated by recent work finding that more than one in five young men hold a positive view of Andrew Tate, a self-proclaimed misogynist who has posted content advocating for the hitting and choking of women (Booth,, 2024; Kings College London,, 2024).
The evidence described above supports the notion that the online space reflects structural inequalities offline, aligning with feminist accounts of how misogyny adapts across contexts, with online spaces now continuations of gendered social orders rather than neutral arenas (e.g., Banet-Weiser and Miltner,, 2016; Lewis et al.,, 2017; Nadim and Fladmore,, 2019). Indeed, further work also suggests that online violence against women often comprises ‘silencing strategies’ such as rape and death threats, attempting to prevent women from partaking in online discussion (Lumsden and Morgan,, 2017). Growing concerns over the problem in recent years has led to the development of the UK Government’s strategy to tackle violence against women and girls in 2021 (Home Office,, 2021), and the recently passed Online Safety Act now requires Ofcom to develop guidance for tech companies to reduce harm to women and girls (End Violence Against Women,, 2023).
@@
However, there is some conflicting evidence regarding gender differences in absolute exposure to online harms, with some recent work finding no differences in overall exposure (Enock et al.,, 2023; Ofcom,, 2023), and some reports even finding that men experience certain harms such as ‘sextortion’ (online blackmail where criminals threaten to release sexual or indecent images unless money is paid) and physical threats to a greater extent than women (BBC,, 2023; Pew Research Center,, 2021). Explaining this apparent contradiction, one possibility is that much work examining online violence against women in the current literature does not include men in the sample, making it difficult to directly measure how gender interacts with nuanced experiences of online harms and perhaps emphasising the experiences of women (e.g., Amnesty International,, 2017; European Women’s Lobby,, 2017). Another possibility is that the kinds of harms affecting women online are qualitatively different to those affecting men, with women more at risk of contact-based harms such as sexual harassment and cyberflashing (Levrant Miceli et al.,, 2001; Gillett,, 2023; European Women’s Lobby,, 2017), meaning differences may go undetected when asking about exposure to harms as a whole. In this work, we aim to deepen our understanding of people’s key experiences, concerns and behaviours surrounding online harms and online safety with a particular focus on when and how women are disproportionately affected. We examine the impact of gender, exposure to harm and fears surrounding such exposure on online behaviours in a nationally representative sample of adults living in the UK.
@@ 1.3 Psychological impacts of online experiences
Beyond gendered inequalities in exposure to online harms, it is also important to consider their psychological impact. Research has found that exposure to online harms can cause severe harm to the psychological wellbeing of those exposed. One large cross-sectional study of Americans and Finns aged 15-30 found that exposure to harm-advocating content such as eating disorder promotion content, self-harm encouragement and sites portraying deaths of others positively was associated with significantly lower subjective wellbeing (Keipi et al.,, 2017). Further, a systematic review found that victims of cyberstalking and harassment commonly reported experiencing depression, fear, anger, suicidal ideation, shame, paranoia and isolation, amongst other negative emotions (Stevens et al.,, 2021). As well as potentially facing greater targeting of certain online harms, women may also suffer more psychologically as a result of such exposure because of the way these abuses operate as forms of ‘symbolic violence’ (described as the normalised ways in which inequality is reproduced and accepted Bourdieu, (1991)). Through this lens, gendered abuse online is not only damaging to individuals but also reinforces women’s unequal position within a patriarchal society. Supporting this, research by Amnesty found that of the 20% of women who had experienced abuse or harassment through social media, 55% suffered stress, anxiety or panic attacks and 61% had trouble sleeping (Amnesty International,, 2017). Additionally, only 3% of women who had experienced abuse or harassment say they were untroubled by these experiences (compared to 17% of men), while 49% of women (compared to 35% of men) stated that their experiences of online abuse made them feel ashamed (Storry and Poppleton,, 2022). Recent work from Ofcom showed that women were less likely than men to say that being online has a positive effect on their mental health (40% of women compared to 45% of men) (Ofcom,, 2022). While some prior work in journalism and media studies has examined the negative psychological impact of online abuse on women (e.g. Chen et al.,, 2020), few studies have systematically compared the psychological experiences of men and women in response to specific online experiences in a representative sample. To address this gap, we examine whether women are more likely than men to be negatively emotionally affected by online experiences.
@@ 1.4 Fears surrounding online experiences
Closely tied to negative psychological experience is heightened fear. Offline, women consistently report higher levels of fear of being targeted by crime compared with men, especially when considering crime relating to personal harm (Snedker,, 2012). Stanko, (1995) argues that women’s concerns around danger are largely caused by a fear of men, reflecting women’s positions in a gendered world. It is important to understand feelings of fear around safety online as these are likely to impact people’s comfort with participating in the online space. While some work suggests that women’s fears around personal safety are heightened compared to men’s, few previous studies have systematically examined gender differences in fears of specific types of online harm, which is a key aim of the present study. We also explore where fears of online harms typically originate from, an important first step in understanding the function of cultural transmission in discussions about online safety.
1.5 Online safety work and participation
Fears surrounding online harms can be understood as part of a broader pattern of gendered safety work, where the burden falls upon women to manage their own risks (Vera-Gray and Kelly,, 2020). In an online context, such safety work often requires engaging with platform safety technologies, such as blocking, reporting, or altering privacy settings (Bright et al.,, 2024). However, little research has explored whether there are gender differences in uptake of such tools, which would align with theoretical claims that women engage in disproportionate self-protection work online. We examine whether women are more likely than men to engage in safety work online, reflected in increased use of safety tools.
@@ 1.5 Online safety work and participation
Offline, fear of crime regularly leads to changes in behaviour. As fear of crime is often higher in women, the restrictive behavioural consequences can be greater for women too (Stanko,, 1990; Snedker,, 2012). This can impact women in a myriad of ways, for example in the extra measures they take to ensure they get home safely at night and additional precautions when travelling alone (Office for National Statistics,, 2022). Just as women work harder to protect their safety offline, it looks increasingly likely that this additional burden must be carried into the online world too. Work from Ofcom (2022) found that women are significantly less likely than men to feel that they are able to have a voice online, while research exploring academics’ interactions with social media found that fears of online harassment in female and ethnic minority academics often leads to self-censorship (Olson Carter and LaPoe,, 2018). These findings show the silencing effect that online abuse can have, with such harms directly threatening freedom of expression. In the digital context, this extra work can place uneven demands on women and sustain digital inequalities (Vera-Gray,, 2018). We investigate the impact of gender on comfort with online participation, such as expressing political opinions and sharing content online, and test whether fears of online harms are associated with reduced comfort with online participation.
2 Research contribution
@@
We seek to address the gaps in the research landscape that remain surrounding when and how women are disproportionately affected by harmful content online. We build on feminist accounts that frame online harms as part of broader patterns of gendered inequality, and ask when and how such harms affect women differently to men.
@@
While much existing work examining online violence against women and girls includes only women and girls in the sample as the key subjects of interest, this means it can be difficult to understand the true extent to which women are affected differently to men. We survey a sample of UK adults nationally representative across age, gender and ethnicity so that we can make meaningful gender comparisons across our variables of interest. Additionally, while much of the existing research either asks about exposure to harmful content in general, or focuses on very specific harms such as abuse and harassment, we ask about exposure and fears relating to a comprehensive list of fifteen potential harms, including content-based ones (those that arise from viewing certain pieces of content) like misinformation and eating disorder promotion content, along with contact-based ones (those that originate from direct behaviours towards the target) such as cyberflashing, cyberstalking and image-based sexual abuse. Importantly, we ask not only about general exposure to such harms, but also about whether people have been directly targeted by each one and to what extent. In doing so, we are able to capture a more detailed picture of the different ways in which men and women experience different types of harmful content online.
@@
Our survey design also allows us to deepen our understanding of when and how women are affected differently by online harms to men. Much previous work in the area looks only at how feelings and behaviours change after harm exposure, and only in samples that report having been affected. In our sample, we are able to compare psychological impacts of online experiences between men and women directly, along with fears, the origins of such fears, and subsequent behavioural changes. We are also able to understand whether women engage in more work overall to protect their safety online, including censoring themselves or behaving more cautiously. We also examine associations between gender, fears and online participation.
3 Methodology
3.1 Data collection and ethics
@@
We conducted a nationally representative survey of 2,000 UK adults. Data collection took place online during June 2023, the survey was created and administered using Qualtrics111www.qualtrics.com and participants were recruited through Prolific222www.prolific.com.The survey was approved by the Ethics Committee at The Alan Turing Institute, UK (approval number C2105-074). Informed consent was obtained at the start of the survey according to approved ethical procedures. The materials and data are available on request.
@@ 3.2 Sample
A total of 1992 participants who completed the survey passed standard checks for data quality and were included in the final sample. The sample was designed to be nationally representative of the population of the United Kingdom across demographic variables of age, gender and ethnicity using Prolific’s representative sample tool. Respondents were between 18 and 90 years old, with a mean age of 45.8 (SD = 15.5). A total of 1010 participants identified as female (50.7%), 965 as male (48.4%), with 9 as non-binary, 5 selecting ‘prefer not to say’, and 3 self-describing. The majority of respondents identified as White (86.9%), while 7.2% as Asian or Asian British, 3.0% as Black, African, Caribbean, or Black British, and 1.6% as mixed, multiple or other ethnicities (0.7% participants selected ‘any other ethnic group’, while 0.6% chose ‘prefer not to say’). [^3] Although participants indicated more specific ethnic identities, we have combined them into broader categories to simplify reporting here. As the main aim of our survey is understanding gender differences, we exclude the 17 participants choosing gender as ’non-binary’, ’prefer not to say’ or ’prefer to self-describe’ from these analyses because the sample size is too small to make meaningful comparisons and inferences for these sub-samples.